A person is pushing a wheelbarrow along a ramp that makes an angle α = 41.0° with the horizontal. The wheelbarrow and load have a combined mass of 21.35 kg with the center of mass at the midpoint of the length L. What is the total force the person must apply so that the wheelbarrow remains horizontal on that ramp?
Answer #1

net force needed to make the wheel barrow horizontal, wben the center of gravity ar mid point of the system

considering moment about wheel

F cos x* L = mg * ( L/2)

F cos 41 = 21.35* 9.8 / 2

F = 138.616 N

========

> Thank you so much after 5 tries and 3 wrong suggested answers, you finally got it correct!

Joshua Vas Fri, Dec 3, 2021 5:49 PM